Meaghan notices that at 6 o'clock the minute hand is on the 12. What is the degree of rotation of the minute hand between 6:00 and 6:20?
Thanks!

Answer by jim_thompson5910(35256)

20 minutes is 20/60 = 1/3 of an hour

So the arc is (1/3)*(360) = 120 degrees

The degree of rotation is 120 degrees